OpenStudy (anonymous):

f(x) = 1/x+2, what is f(3+h)?

OpenStudy (psymon):

f(________) whatever goes in that blank will replace every single x in the function with that blank. So for example, if you have f(x) = 2x^2 -x, then f(3) = 2(3)^2 - 3 f(x+5) = 2(x+5)^2 - (x+5) f(something) = 2(something)^2 - (something) So its the same thing with your problem, it just means replace x with 3 + h

OpenStudy (agent0smith):

Replace x with 3+h, and you're done.

OpenStudy (anonymous):

So it would be 1/3+h

OpenStudy (anonymous):

1/5+h

OpenStudy (psymon):

You only replace x with 3 + h, it doesnt make the 2 disappear suddenly.

OpenStudy (anonymous):

Yeah I meant 1/5+h right?

OpenStudy (psymon):

yeah